Review Constants Part A A Doppler blood flow unit emits ultrasound at 5.2 MHz What is the frequency shift of the ultrasound reflected from blood moving in an artery at a speed of 0.18 Express your answer to two significant figures and include the appropriate units. m/s? df= 1 Value Units Submit
Answer #1

Solution)

Here, frequency f= fo(v+vb/v-vb)

We know, speed of sound in blood is 1540 m/s

F= 5.2*10^6(1540+ 0.18/1540-0.18)

F= 5.201215 Mhz

Frequency shift, delta F= F-Fo

Delta F= (5.201215 )- (5.2)=0.001215 MHz = 1.21 * 10^3 Hz (Ans)

=====